A beaker contains 600 ml of a 8% saltwater solution. How much water must be boiled away in order to have a 9% solution? Enter your answer correct to two decimal places.
Answer by Alan3354(69443)

A beaker contains 600 ml of a 8% saltwater solution. How much water must be boiled away in order to have a 9% solution? Enter your answer correct to two decimal places.
---------------
600*0.08 = 48 ml of salt
----
If 48 is 9% = 0.09:
Total solution = 48/0.09 =~533.33 ml
---
600 - 533.33 = 66.67 to boil away.